OpenStudy (anonymous):

The point-slope form of the equation of the line that passes through (–5, –1) and (10, –7) is y + 7 -2/5 = (x – 10). What is the standard form of the equation for this line? A. 2x – 5y = –15 B. 2x – 5y = –17 C. 2x + 5y = –15 D. 2x + 5y = –17

OpenStudy (dayakar):

equation u written is wrong plz check a small correction is needed

OpenStudy (faiqraees):

C is the correct answer. Just take any point like (10,-7) and plug it into every equation. If the equation is correct, the right-hand side and the left-hand side of the equation will be equal. For C 2x+5y=-15 X=10 y=-7 2(10)+5(-7)=-15 20-35=-15 -15=-15